the cost of a new car included a 4 % sales tax. if c represents the cost of the car before tax, write an expression for the total cost, including sales tax.

Answer by Theo(13342)

c = cost of car before taxes.
c + .04 * c = cost of car after taxes.
you can factor out the c to get:
c * (1 + .04) = cost of car after taxes.
you can simplify to get:
c * 1.04 = cost of car after taxes.
this can also be written as:
1.04 * c = cost of car after taxes.